Reid Company is considering the production of a new product. The expected variable cost is $28 per unit. Annual fixed costs are expected to be $630,000. The anticipated sales price is $70 each. Required Determine the break-even point in units and dollars using each of the following: Use the equation method. Use the contribution margin per unit approach. Use the contribution margin ratio approach. (Do not round intermediate calculations. Round "Contribution margin ratio" to 1 decimal place. (i.e., 0.234 should be entered as 23.4))
Answer #1

contribution margin per unit approach.

Sale per unit - Variable cost per unit = contribution margin per unit

70 - 28 = 42

42 = contribution margin per unit

contribution margin ratio approach.

contribution margin ratio = Fixed Cost / Contribution margin

Contribution margin = Sale - variable cost / sales

Contribution margin = 70 - 28 / 70

Contribution margin =60%

contribution margin ratio = 630000 / 0.60 = 1050000

Break even points in units = Fixed cost / Sale per unit - Variable cost per unit

Break even points in units = 630000 / 70-28

Break even points in units = 630000 / 42

Break even points in units =15000 units

Break even points in dollars = Fixed Cost / Contribution margin

Contribution margin = Sale - variable cost / sales

Contribution margin = 70 - 28 / 70

Contribution margin =60%

contribution margin ratio = 630000 / 0.60 = 1050000